# CAGE: Continuity-Aware edGE Network Unlocks Robust Floorplan Reconstruction

### v1

@@ Abstract
We present CAGE (Continuity-Aware edGE) network, an end-to-end framework for reconstructing vector floorplans directly from point-cloud density maps. Traditional corner-based polygon representations are highly sensitive to noise and incomplete observations, often resulting in fragmented or implausible layouts. Recent line grouping methods leverage structural cues to improve robustness but still struggle to recover fine geometric details. To address these limitations, we propose a native edge-centric formulation, modeling each wall segment as a directed, geometrically continuous edge. This representation enables inference of coherent floorplan structures, ensuring watertight, topologically valid room boundaries while improving robustness and reducing artifacts. Towards this design, we develop a dual-query transformer decoder that integrates perturbed and latent queries within a denoising framework, which not only stabilizes optimization but also accelerates convergence. Extensive experiments on Structured3D and SceneCAD show that CAGE achieves state-of-the-art performance, with F1 scores of 99.1% (rooms), 91.7% (corners), and 89.3% (angles). The method also demonstrates strong cross-dataset generalization, underscoring the efficacy of our architectural innovations. Code and pretrained models will be released upon acceptance.
@@ 1 Introduction
Reconstructing indoor scenes into compact, editable vector floorplans is a longstanding goal in computer vision and robotics [1, 2]. A vector floorplan is a structured 2D representation of interior geometry, composed of lines or polygons that delineate walls, room boundaries, and architectural elements. Unlike raster maps, vector floorplans are resolution-independent, support precise geometric reasoning, and integrate seamlessly with CAD tools and building information models (BIM). Their joint encoding of spatial topology and geometry makes them ideal for downstream applications such as building lifecycle management, AR/VR simulation, and autonomous navigation [3, 4, 5, 6].
@@
Modern floorplan reconstruction uses diverse inputs including RGB images [7], panoramic views [8], CAD drawings [9], and 3D point clouds [10]. Among these, projecting point clouds to 2D density maps provides an optimal balance of geometric accuracy and computational efficiency [11, 12], making it popular in learning-based approaches. However, real-world scans often suffer from severe occlusions, clutter, and incompleteness that obscure critical structural elements. Additionally, the discretization process in density map creation blurs corner features and wall segments. These challenges significantly complicate accurate vector floorplan recovery. The resulting incomplete or noisy representations demand robust algorithms capable of handling such imperfect data.
@@
Notably, a critical factor affecting reconstruction quality lies in the choice of floorplan representation. Recent work has explored a range of methods. For example, HEAT [13] detects discrete corners and connects them into edges and polygons, yet it often suffers from gaps caused by missed detections. RoomFormer [12] models floorplans as sequences of corners using a two-level query transformer, yet a single missing corner can distort the entire layout. As a result, corner-based methods are highly sensitive to noise. SLIBO-Net [14] introduces a slicing-box representation but relies heavily on the Manhattan-world assumption. FRI-Net [15] reconstructs rooms by compositional line grouping, its implicit neural representations often lead to over-smoothing of fine structures. It is clear current forms still struggle to capture global structure while keeping local precision.
@@
To address these challenges, we propose CAGE (Continuity-Aware edGE) network, an end-to-end framework for reconstructing vector floorplans from density maps. Our CAGE adopts an edge-centric formulation, modeling each wall as a directed, geometrically continuous edge. This design improves robustness to incomplete or noisy data by removing reliance on precise corner localization, while enabling global reasoning. To complement this representation, we introduce a dual-query transformer decoder that integrates perturbed and latent edge queries within a denoising framework, enhancing training stability and accelerating convergence. As decoding progresses, edge predictions are iteratively refined, yielding clean and watertight polygons. Evaluations on Structured3D [16] and SceneCAD [17] demonstrate that CAGE achieves state-of-the-art performance and generalizes well across datasets. Our main contributions are as follows:
@@
  • Dual-Query Transformer Decoder: Our approach leverages perturbed and latent queries within a denoising framework to stabilize training, refine predictions, accelerate convergence.
  • Strong Performance and Generalization: We achieve state-of-the-art performance on multiple benchmarks and generalizes effectively across datasets, validating the effectiveness of our continuity-aware edge based formulation.
@@ 2.1 Traditional Methods
Early work in floor- and room-layout recovery relied on low-level vision cues or geometric heuristics. Some methods extract planes from point clouds and optimize line placements to assemble floorplans [18, 19], while others infer structure from panoramic images [20, 8] or CAD drawings [9, 21, 22]. Techniques such as RANSAC-based plane fitting and piecewise merging have been used to generate compact surface meshes that are later flattened into floorplans [23]. Graph-based formulations often cast the reconstruction as shortest-path or global optimization problems, producing watertight layouts but requiring hand-designed potentials [24, 25]. More recent pipelines focus on handling cluttered LiDAR scans via global optimization over curved surfaces [26] or decomposing scenes into walls and objects for lightweight BIM generation [27]. From a single panorama, Pano2CAD estimates Manhattan room geometry and object pose [28], while successors like PanoFormer [29] support curved wall recovery via tessellated spherical surfaces. Recent work also explores LoD4 building modeling by fusing interior and exterior cues from SfM images [30]. Traditional methods require small amounts of training data while heavily rely on the design of handcraft features.
@@ 2.2 Stage-wise Learning-based Methods
Several methods adopt a hybrid pipeline where learning components are combined with post-processing optimization. FloorNet [31] detects corners in a top-view raster and assembles them via integer programming, while HEAT [13] replaces the solver with a transformer-based corner-pair classifier. Floor-SP [11] and MonteFloor [32] segment rooms using Mask-RCNN or FCN, refine initial contours through shortest path or Monte Carlo Tree Search, and encourage geometric consistency by sharing corners and walls. PolyGraph [36] generates wall points using a cross-guided neural network, forms initial triangles, and applies post-processing to refine polygonal layouts. FloorUSG [19] integrates 2D plane instances with 3D geometry to lift RGB features into structured floorplans. ArrangementNet [10] detects walls and partitions space using wall lines, then leverages an extended GCN to model collinearity and coplanarity among surface patches, improving segmentation in complex scenes. SLIBO-Net [14] introduces a slicing box representation with geometric regularization and post-processing for capturing fine local details. While these stage-wise systems achieve strong performance, they are often sensitive to missing primitives and rely on extensive post-optimization.
@@ 2.3 End-to-End Learning-based Methods
Recent advances favor end-to-end pipelines that directly predict vector floorplan using transformer-based decoders. RoomFormer [12] introduces a transformer architecture with hierarchical queries to jointly predict multiple corner sequences. Although effective, it suffers from unordered predictions due to random query initialization and incomplete polygons caused by missed corners. PolyRoom [37] enhances this framework by segmenting rooms with Mask-RCNN and initializing RoomFormer-style queries from detected contours. PolyDiffuse [38] formulates floorplan generation as a conditional task and introduces a guided set diffusion model for room representation optimization. These corner-based methods treat floorplan reconstruction as discrete corner regression, often neglecting global shape coherence. FRI-Net [15] addresses this by learning room-wise latent codes and decoding them into lines, which are grouped into polygons using a BSP-Net-inspired [39] grouping strategy. Recent methods have progressed from corner-based to line-grouping approaches, hightlighting the importance of structural information. However, they still fail to balance robustness with geometric precision. We address this with edge-centric representation that fundamentally reformulates the problem, enabling accurate, topologically valid reconstructions in an end-to-end framework.
@@ 3 Edge-based Floorplan Representation
We propose an explicit edge-based representation for modeling floorplan, addressing the key limitations of prior methods, see Figure 2(a)(b). We represent each room polygon as an ordered sequence of directed edges, where each edge is defined by a pair of endpoints in normalized 2D space, illustrated in Figure 2(b). Let a floorplan contain at most $M$ rooms, and each room contain at most $N$ edges. We represent the $m$-th room as:
@@
$T_{m}=\{t_{m}^{n}\}_{n=1}^{N},\quad\text{where }N\geq\max_{m}N_{m}$ (3)
@@ 4.1 Overview
Figure 4 illustrates the overall architecture of our proposed CAGE framework for end-to-end floorplan reconstruction from 3D point clouds. The input point cloud is first projected onto the XY-plane to generate a 2D density map, which is processed by a convolutional image backbone to extract multi-scale visual features. The output feature maps are flattened and augmented with positional encodings to form a unified token sequence, which is passed to a transformer encoder. We employ multi-scale deformable attention [40] in the encoder to efficiently aggregate both local and global spatial context across all feature levels.
The encoded features are processed by our novel transformer decoder composed of stacked attention layers. The decoder operates on a set of learnable edge queries, which are refined across layers through self-/cross-attention mechanisms. Each query predicts a directed edge denoted by two endpoints and is assigned a binary label indicating whether it constitutes a valid segment, as described in Sec. 3. To improve the robustness and training stability, we incorporate a dual-query design that includes latent queries for final predictions and perturbed queries for denoising supervision. The following subsections will illustrate our novel dual-query decoding mechanism, and loss formulation.
@@ 4.2 Iterative Polygon Refinement by Dual Query
As described in Sec. 3, we represent each room as a sequence of edges and flatten all edges into query tensor as $Q\in\mathbb{R}^{M\times N\times 2}$, where $M$ is the maximum number of polygons and $N$ the number of edges per polygon. Each query predicts an edge defined by two endpoints. The decoder layers refine these queries iteratively, as illustrated in Figure 5, where edge geometry and corresponding polygons become increasingly accurate across decoding steps.
@@
Each perturbed edge query is generated by applying controlled noise to a ground-truth edge during training. Specifically, for each edge, its two endpoints are randomly displaced to simulate input uncertainty. The displacements are constrained relative to the spatial extent of the density map: the horizontal shift $\Delta x$ and vertical shift $\Delta y$ are bounded by $|\Delta x|<\lambda w/2$ and $|\Delta y|<\lambda h/2$, where $w$ and $h$ denote the width and height of the input density image, respectively. If a perturbed endpoint falls outside the image boundary, it is clamped to remain within the valid range. The scalars $\lambda$ is hyperparameter that determine the magnitude of perturbation. Additionally, to introduce label noise, the binary class label of each edge is randomly flipped with probability $\gamma$. This denoising strategy encourages the decoder to recover clean edge structures from noisy inputs. During inference, only the latent queries are used to generate final polygons.
@@ 4.3 Loss Functions
Polygon Matching. Let the predicted edge set be $\hat{S}=\{\hat{E}_{m}=(\hat{e}_{m}^{1},...,\hat{e}_{m}^{N})\}_{m=1}^{M}$, where each edge $\hat{e}_{m}^{n}=(\hat{c}_{m}^{n},\hat{p}_{m1}^{n},\hat{p}_{m2}^{n})$ includes a binary confidence score and two endpoint coordinates. Ground-truth polygons are denoted as $S=\{E_{m}\}_{m=1}^{M_{\text{gt}}}$, each padded to length $N$ and further extended with dummy polygons for matching. We compute a bipartite matching between predicted and ground-truth polygons using the Hungarian algorithm, minimizing the following assignment cost:
@@
Loss Components. After matching, we supervise the predicted polygons using three terms: classification, edge regression, and rasterization. For latent queries, losses are applied only to matched predictions. For perturbed queries, losses are computed without matching.
@@ 5.1 Experiments Setting
Dataset and Evaluation Metrics. We evaluate our method on two large-scale indoor datasets, including Structured3D [16] and SceneCAD [17]. Structured3D is a photo-realistic synthetic dataset comprising 3,500 houses with diverse Manhattan and non-Manhattan layouts. SceneCAD provides 3D room layout annotations for real-world RGB-D scans from ScanNet [44], with each sample containing a single room. Following previous work [13, 12], we split Structured3D into 3,000 training, 250 validation, and 250 test samples. For SceneCAD, we use the provided splits of 828 training and 127 validation samples. Following [32, 13], we process registered multi-view RGB-D panoramas into point clouds, and project them vertically into $256\times 256$ pixel density images. Each pixel value is normalized to $[0,1]$ by dividing the number of projected points by the maximum count per image. In line with prior work [13, 12, 36], we evaluate performance using Precision, Recall, and F1 scores at the Room, Corner, and Angle levels.
Baselines. We compare our method with eight state-of-the-art approaches: Floor-SP [11], HEAT [13], RoomFormer [12], SLIBO-Net [14], FRI-Net [15], PolyRoom [37], PolyGraph [36], and PolyDiffuse [38]. Floor-SP [11] uses Mask-RCNN for room segmentation followed by classical optimization for vectorization. HEAT [13] detects corners and infers connectivity with a neural network. RoomFormer [12] employs a transformer to jointly predict multiple corner sequences. SLIBO-Net [14] generates slicing boxes and room centroids via transformer decoding, refined through post-processing. FRI-Net [15] encodes room-wise latent features, decodes them into lines, and assembles polygons using a BSP-Net-inspired [39] grouping strategy. PolyRoom [37] combines Mask-RCNN-based segmentation with RoomFormer-style reconstruction initialized from detected contours. PolyGraph [36] predicts wall points and triangulations and then refines results by postprocessing. PolyDiffuse [38] formulates reconstruction as a conditional generation task, using a diffusion model to iteratively refine floorplans from initial proposals.
@@
Our model is implemented in PyTorch and trained on a single NVIDIA RTX 4090 GPU. We use Swin Transformer V2 [45] as the default image backbone, while ResNet-50 [46] and Swin Transformer V1 [47] are used in ablation studies. The transformer architecture follows a standard 6-layer encoder-decoder configuration. We apply denoising training with dual-query supervision and optimize using Adam. Further model configurations, training schedules, and hyperparameters are provided in supplementary material A.

### v2

@@ Abstract
We present CAGE (Continuity-Aware edGE) network, a robust framework for reconstructing vector floorplans directly from point-cloud density maps. Traditional corner-based polygon representations are highly sensitive to noise and incomplete observations, often resulting in fragmented or implausible layouts. Recent line grouping methods leverage structural cues to improve robustness but still struggle to recover fine geometric details. To address these limitations, we propose a native edge-centric formulation, modeling each wall segment as a directed, geometrically continuous edge. This representation enables inference of coherent floorplan structures, ensuring watertight, topologically valid room boundaries while improving robustness and reducing artifacts. Towards this design, we develop a dual-query transformer decoder that integrates perturbed and latent queries within a denoising framework, which not only stabilizes optimization but also accelerates convergence. Extensive experiments on Structured3D and SceneCAD show that CAGE achieves state-of-the-art performance, with F1 scores of 99.1% (rooms), 91.7% (corners), and 89.3% (angles). The method also demonstrates strong cross-dataset generalization, underscoring the efficacy of our architectural innovations. Code and pretrained models are available on our project page: https://github.com/ee-Liu/CAGE.git.
@@ 1 Introduction
Reconstructing indoor scenes into compact, editable vector floorplans is a longstanding goal in computer vision and robotics [1; 2]. A vector floorplan is a structured 2D representation of interior geometry, composed of lines or polygons that delineate walls, room boundaries, and architectural elements. Unlike raster maps, vector floorplans are resolution-independent, support precise geometric reasoning, and integrate seamlessly with CAD tools and building information models (BIM). Their joint encoding of spatial topology and geometry makes them ideal for downstream applications such as building lifecycle management, AR/VR simulation, and autonomous navigation [3; 4; 5; 6].
@@
Modern floorplan reconstruction uses diverse inputs including RGB images [7], panoramic views [8], CAD drawings [9], and 3D point clouds [10]. Among these, projecting point clouds to 2D density maps provides an optimal balance of geometric accuracy and computational efficiency [11; 12], making it popular in learning-based approaches. However, real-world scans often suffer from severe occlusions, clutter, and incompleteness that obscure critical structural elements. Additionally, the discretization process in density map creation blurs corner features and wall segments. These challenges significantly complicate accurate vector floorplan recovery. The resulting incomplete or noisy representations demand robust algorithms capable of handling such imperfect data.
@@
Notably, a critical factor affecting reconstruction quality lies in the choice of floorplan representation. Recent work has explored a range of methods. For example, HEAT [13] detects discrete corners and connects them into edges and polygons, yet it often suffers from gaps caused by missed detections. RoomFormer [12] models floorplans as sequences of corners using a two-level query transformer, yet a single missing corner can distort the entire layout. As a result, corner-based methods are highly sensitive to noise. SLIBO-Net [14] introduces a slicing-box representation but relies heavily on the Manhattan-world assumption. FRI-Net [15] reconstructs rooms by compositional line grouping, its implicit neural representations often lead to over-smoothing of fine structures. It is clear that current forms still struggle to capture global structure while keeping local precision.
@@
To address these challenges, we propose CAGE (Continuity-Aware edGE) network, a robust framework for reconstructing vector floorplans from density maps. Our CAGE adopts an edge-centric formulation, modeling each wall as a directed, geometrically continuous edge. This design improves robustness to incomplete or noisy data by removing reliance on precise corner localization, while enabling global reasoning. To complement this representation, we introduce a dual-query transformer decoder that integrates perturbed and latent edge queries within a denoising framework, enhancing training stability and accelerating convergence. As decoding progresses, edge predictions are iteratively refined, yielding clean and watertight polygons. Evaluations on Structured3D [16] and SceneCAD [17] demonstrate that CAGE achieves state-of-the-art performance and generalizes well across datasets. Our main contributions are as follows:
@@
  • Dual-Query Transformer Decoder: Our approach leverages perturbed and latent queries within a denoising framework to stabilize training, refine predictions, and accelerate convergence.
  • Strong Performance and Generalization: We achieve state-of-the-art performance on multiple benchmarks and generalize effectively across datasets, validating the effectiveness of our continuity-aware edge-based formulation.
@@ 2.1 Traditional Methods
Early work in floor- and room-layout recovery relied on low-level vision cues or geometric heuristics. Some methods extract planes from point clouds and optimize line placements to assemble floorplans [18; 19], while others infer structure from panoramic images [20; 8] or CAD drawings [9; 21; 22]. Techniques such as RANSAC-based plane fitting and piecewise merging have been used to generate compact surface meshes that are later flattened into floorplans [23]. Graph-based formulations often cast the reconstruction as shortest-path or global optimization problems, producing watertight layouts but requiring hand-designed potentials [24; 25]. More recent pipelines focus on handling cluttered LiDAR scans via global optimization over curved surfaces [26] or decomposing scenes into walls and objects for lightweight BIM generation [27]. From a single panorama, Pano2CAD estimates Manhattan room geometry and object pose [28], while successors like PanoFormer [29] support curved wall recovery via tessellated spherical surfaces. Recent work also explores LoD4 building modeling by fusing interior and exterior cues from SfM images [30]. Traditional methods require small amounts of training data while heavily relying on the design of handcrafted features.
@@ 2.2 Stage-wise Learning-based Methods
Several methods adopt a hybrid pipeline where learning components are combined with post-processing optimization. FloorNet [31] detects corners in a top-view raster and assembles them via integer programming, while HEAT [13] replaces the solver with a transformer-based corner-pair classifier. Floor-SP [11] and MonteFloor [32] segment rooms using Mask-RCNN [33], refine initial contours through shortest path or Monte Carlo Tree Search [34; 35], and encourage geometric consistency by sharing corners and walls. PolyGraph [36] generates wall points using a cross-guided neural network, forms initial triangles, and applies post-processing to refine polygonal layouts. FloorUSG [19] integrates 2D plane instances with 3D geometry to lift RGB features into structured floorplans. ArrangementNet [10] detects walls and partitions space using wall lines, then leverages an extended GCN to model collinearity and coplanarity among surface patches, improving segmentation in complex scenes. SLIBO-Net [14] introduces a slicing box representation with geometric regularization and post-processing for capturing fine local details. While these stage-wise systems achieve strong performance, they are often sensitive to missing primitives and rely on extensive post-optimization.
@@ 2.3 End-to-End Learning-based Methods
Recent advances favor end-to-end pipelines that directly predict vector floorplan using transformer-based decoders. RoomFormer [12] introduces a transformer architecture with hierarchical queries to jointly predict multiple corner sequences. Although effective, it suffers from unordered predictions due to random query initialization and incomplete polygons caused by missed corners. PolyRoom [37] enhances this framework by segmenting rooms with Mask-RCNN [33] and initializing RoomFormer-style queries from detected contours. PolyDiffuse [38] formulates floorplan generation as a conditional task and introduces a guided set diffusion model for room representation optimization. These corner-based methods treat floorplan reconstruction as discrete corner regression, often neglecting global shape coherence. FRI-Net [15] addresses this by learning room-wise latent codes and decoding them into lines, which are grouped into polygons using a BSP-Net-inspired [39] grouping strategy. Recent methods have progressed from corner-based to line-grouping approaches, highlighting the importance of structural information. However, they still fail to balance robustness with geometric precision. We address this with edge-centric representation that fundamentally reformulates the problem, enabling accurate, topologically valid reconstructions in an end-to-end trainable framework.
@@ 3 Edge-based Floorplan Representation
We propose an explicit edge-based representation for modeling floorplan, addressing the key limitations of prior methods, see Figure 2(a)(b). We represent each room polygon as an ordered sequence of directed edges, where each edge is defined by a pair of endpoints in normalized 2D space, as illustrated in Figure 2(b). Let a floorplan contain at most $M$ rooms, and each room contain at most $N$ edges. We represent the $m$-th room as:
@@
$T_{m}=\{t_{m}^{n}\}_{n=1}^{N},\quad\text{where }\max_{m}N_{m}\leq N$ (3)
@@ 4.1 Overview
Figure 4 illustrates the overall architecture of our proposed CAGE framework for floorplan reconstruction from 3D point clouds. The input point cloud is first projected onto the XY-plane to generate a 2D density map, which is processed by a convolutional image backbone to extract multi-scale visual features. The output feature maps are flattened and augmented with positional encodings to form a unified token sequence, which is passed to a transformer encoder. We employ multi-scale deformable attention [40] in the encoder to efficiently aggregate both local and global spatial context across all feature levels.
The encoded features are processed by our novel transformer decoder composed of stacked attention layers. The decoder operates on a set of learnable edge queries, which are refined across layers through self-/cross-attention mechanisms. Each query predicts a directed edge denoted by two endpoints and is assigned a binary label indicating whether it constitutes a valid segment, as described in Sec. 3. To improve the robustness and training stability, we incorporate a dual-query design that includes latent queries for final predictions and perturbed queries for denoising supervision. The following subsections will illustrate our novel dual-query decoding mechanism and loss formulation.
@@ 4.2 Iterative Polygon Refinement by Dual Query
As described in Sec. 3, we represent each room as a sequence of edges and flatten all edges into query tensor as $Q\in\mathbb{R}^{M\times N\times 2}$, where $M$ is the maximum number of polygons and $N$ is the maximum number of edges per polygon. Each query predicts an edge defined by two endpoints. The decoder layers refine these queries iteratively, as illustrated in Figure 5, where edge geometry and corresponding polygons become increasingly accurate across decoding steps.
@@
Each perturbed edge query is generated by applying controlled noise to a ground-truth edge during training. Specifically, for each edge, its two endpoints are randomly displaced to simulate input uncertainty. The displacements are constrained relative to the spatial extent of the density map: the horizontal shift $\Delta x$ and vertical shift $\Delta y$ are bounded by $|\Delta x|<\lambda w/2$ and $|\Delta y|<\lambda h/2$, where $w$ and $h$ denote the width and height of the input density image, respectively. If a perturbed endpoint falls outside the image boundary, it is clamped to remain within the valid range. The scalar $\lambda$ is a hyperparameter that determines the magnitude of perturbation. Additionally, to introduce label noise, the binary class label of each edge is randomly flipped with probability $\gamma$. This denoising strategy encourages the decoder to recover clean edge structures from noisy inputs. During inference, only the latent queries are used to generate final polygons.
@@ 4.3 Loss Functions
Polygon Matching. Let the predicted edge set be $\hat{S}=\{\hat{E}_{m}=(\hat{e}_{m}^{1},...,\hat{e}_{m}^{N})\}_{m=1}^{M}$, where each edge $\hat{e}_{m}^{n}=(\hat{c}_{m}^{n},\hat{p}_{m1}^{n},\hat{p}_{m2}^{n})$ includes a binary confidence score and two endpoint coordinates. Ground-truth polygons are denoted as $S=\{E_{m}\}_{m=1}^{M_{\text{gt}}}$, each padded to length $N$ and further extended with mock polygons for matching. We compute a bipartite matching between predicted and ground-truth polygons using the Hungarian algorithm, minimizing the following assignment cost:
@@
Loss Components. After matching, we supervise the predicted polygons using three terms: classification, edge regression, and rasterization. For latent queries, losses are applied only to matched predictions. For perturbed queries, losses are directly applied without the matching step, as their correspondence to ground-truth edges is already known.
@@ 5.1 Experiments Setting
Dataset and Evaluation Metrics. We evaluate our method on two large-scale indoor datasets, including Structured3D [16] and SceneCAD [17]. Structured3D is a photo-realistic synthetic dataset comprising 3,500 houses with diverse Manhattan and non-Manhattan layouts. SceneCAD provides 3D room layout annotations for real-world RGB-D scans from ScanNet [44], with each sample containing a single room. Following previous work [13; 12], we split Structured3D into 3,000 training, 250 validation, and 250 test samples. For SceneCAD, we use the provided splits of 828 training and 127 validation samples. Following [32; 13], we process registered multi-view RGB-D panoramas into point clouds, and project them vertically into $256\times 256$ pixel density images. Each pixel value is normalized to $[0,1]$ by dividing the number of projected points by the maximum count per image. In line with prior work [13; 12; 36], we evaluate performance using Precision, Recall, and F1 scores at the Room, Corner, and Angle levels.
Baselines. We compare our method with eight state-of-the-art approaches: Floor-SP [11], HEAT [13], RoomFormer [12], SLIBO-Net [14], FRI-Net [15], PolyRoom [37], PolyGraph [36], and PolyDiffuse [38]. Floor-SP [11] uses Mask-RCNN for room segmentation followed by classical optimization for vectorization. HEAT [13] detects corners and infers connectivity with a neural network. RoomFormer [12] employs a transformer to jointly predict multiple corner sequences. SLIBO-Net [14] generates slicing boxes and room centroids via transformer decoding, refined through post-processing. FRI-Net [15] encodes room-wise latent features, decodes them into lines, and assembles polygons using a BSP-Net-inspired [39] grouping strategy. PolyRoom [37] combines Mask-RCNN-based segmentation with RoomFormer-style reconstruction initialized from detected contours. PolyGraph [36] predicts wall points and triangulations and then refines results by postprocessing. PolyDiffuse [38] formulates reconstruction as a conditional generation task, using a diffusion model to refine floorplans from initial proposals iteratively.
@@
Our model is implemented in PyTorch and trained on a single NVIDIA RTX 4090 GPU. We use Swin Transformer V2 [45] as the default image backbone, while ResNet-50 [46] and Swin Transformer V1 [47] are used in ablation studies. The transformer architecture follows a standard 6-layer encoder-decoder configuration. We apply denoising training with dual-query supervision and optimize using Adam [48]. Further model configurations, training schedules, and hyperparameters are provided in supplementary material A.